On planet Tehar, the acceleration of gravity is the same as that on Earth but there is also a strong downward electric field with the field being uniform close to the planet's surface. A 2.00 kg ball having a charge of 5.01 µC is thrown upward at a speed of 20.1 m/s and it hits the ground after an interval of 4.10 s. What is the potential difference between the starting point and the top point of the trajectory? kV
Answer #1

given

m = 2.00 kg

q = charge of 5.01 µC

= 5.01 x 10-6 C

v = speed of 20.1 m/s

t = interval of 4.10 s

net force is F = E q + m g

F =ma

ma = E q + m g

a = (E q + m g) / m

and y = u t - 1/2 a t2

a = 2 u /t

= 2 x 20.1 / 4.1

a = 9.8048 m/sec2

9.8048 = (E q + m g) / m

(E q + m g) / m = 9.8048

(E x 5.01 x 10-6 + 2 x 9.8) / 2 = 9.8048

(E x 5.01 x 10-6 + 2 x 9.8) = 19.609

E x 5.01 x 10-6 = 9.75 x 10-3

E = 9.75 x 10-3 / 5.01 x 10-6

E = 1946.107 N/C

here T = 4.1 /2

T = 2.05 sec to reach max height

using d = u T - 1/2 a T2

d = 20.1 x 2.05 - 0.5 x 9.8 x 2.052

d = 20.6 m

V = E d

= 1946.107 x 20.6

V = 40089.82036 volts

V = 40.08982036 Kv

so the potential difference between the starting point and the

top point of the trajectory is 40.08982036 KV